Give an example which satisfies the properties of vector space, subspace and inner product space and number should be complex


1
Expert's answer
2021-07-15T10:19:17-0400

Vector space:

A vector space is a set of objects which is called vectors, which may be added together and multiplied together by a numbers, called scalers.

Example:

"V=c_1u_1+c_2u_2+c_3u_3"

Subspace:

A subspace of a vector space V is a subset of H of V which have 3 properties -

  • the zero vector of v is in H
  • H is closed under the vector addition that is for each u and v in H, the sum u+v in H.
  • H is closed under multiplication by scalers that is for each u in H and each scaler c, the vector cu is in H.

Example:

  • u+v=v+u
  • (u+v)+w=u+(v+w)
  • u+(-u)=0

vector space V • {0}

The trivial space {0} is a subspace of V.

Ex. "V = R_2" .

The line x − y = 0 is a subspace of "R_2" .

Vector inner product space :

Let u, v and w be a vector in a vector space V and let c be any scaler. An inner product on V is function that associates a real number <u,v> with each pair of vector u and v and satisfies the following axioms -

<u,v>=<v,u>

<u,v+w>=<u,v>+<u,w>

c<u,v>=<cu,v>

<u,u> and <v,v>=0 if v=0

Example:

"\u3008A, B\u3009= tr(B^TA)"

"A=\begin{bmatrix} 
a_{11} & a_{12} \\
a_{21} & a_{22} \\
a_{31} & a_{32} \\
\end{bmatrix}"


"B=\begin{bmatrix} 
b_{11} & b_{12} \\
b_{21} & b_{22} \\
b_{31} & b_{32} \\
\end{bmatrix}"


"B^TA=A=\begin{bmatrix} 
a_{11} b_{11}+b_{21}a_{21}+b_{31}a_{31}& b_{11} a_{12}+b_{21}a_{22}+b_{31}a_{32} \\
b_{12} a_{11}+b_{22}a_{21}+b_{32}a_{31}& b_{12} a_{12}+b_{22}a_{22}+b_{32}a_{32} \\

\end{bmatrix}"


"<A,B>=b_{11}a_{11}+b_{21}a_{21}+b_{31}a_{31}+b_{12}a_{12}+b_{22}a_{22}+b_{32}a_{32}"


"=\Sigma_{i=1}^{3}\Sigma_{j=1}^{3}a_{ij}b_{ij}"

Hence, we can say that the inner product space "(M_{3,2,<,>})" is isomorphic to the euclidean space "(R^{3\times 2},.)"